Can you help me understand questions like 20s+4=13s-10
Answer by Flannery(124)

What you want to do is isolate s on the right side of the equal sign like this.
20s+4=13s-10
20s-13s+4=13s-13s-10
7s+4=-10
7s+4-4=-10-4
7s=-14
s=-2
Check:
20(-2)+4=13(-2)-10
-40+4=-26-10
-36=-36